6. The emission spectrum of Na(NaNNa [Ne]3s) is a doublet. [8 points) (a) What are the term symbols for excited states?
Answer #1

this is a p1 configuration. For a p electron

l=1and as there is only one electron, L = 1.

For a single electron s = ½ and therefore, S = ½.

Hence, 2S+1 = 2.

J values are:

J = 1 + ½, …… 1- ½ = 3/2, …, ½ = 3/2, ½.

Hence the term symbols are,

2P3/2 and 2P1/3